# Harmonic functions on Tutte’s embeddings and linearized Monge–Ampère equation

### v1

@@ 1.1. Motivation
Convergence of discrete harmonic functions in the small mesh size limit is a classical topic dating back to the first works on numerical methods [Lusternik1926, courant-friedrichs-lewy-1928]. In two dimensions, this question is intrinsically related to the notion of discrete holomorphic functions and their convergence, with important contributions made in the middle of the last century by Ferrand [lelong-ferrand-book, Chapter V] who greatly developed this theory on the square grid and Duffin [duffin-rhombic] who, in particular, suggested a generalization to the so-called rhombic tilings of the complex plane. From the combinatorial perspective, two related classical setups for which the notions of discrete holomorphic and harmonic functions make perfect sense are Brooks–Stone–Smith–Tutte’s square (or, more generally, rectangular) tilings [brooks-smith-stone-tutte-1940] and Tutte’s barycentric, or harmonic, embeddings [tutte1963draw]. One of the sources of the modern interest to discrete complex analysis is the 2d statistical physics; e.g., see [smirnov-icm2010]. In particular, the framework of rhombic lattices was important for the study of scaling limits of free fermionic models with Baxter’s Z-invariant weights; e.g., see [mercat-cmp2001, kenyon-isorad, chelkak-smirnov-universality, boutillier-detiliere-rashel-massiveLaplacian].
@@
Recent advances of the probabilistic approach to the Liouville Quantum Gravity (LQG) in 2d (e.g., see [gwynne-holden-sun-matingLQG-survey] and references therein) motivate a development of discrete complex analysis techniques for possibly very irregular planar grids that appear as embeddings of given (random) planar graphs into the plane. In particular, applications of Tutte’s embeddings and square tilings in the LQG context recently appeared in [gwynne-miller-sheffield-Tutte-mated-CRT, bertacco-gwynne-sheffield-smith-embeddings]. Going in this direction, the last three authors of this paper suggested a general framework of the so-called t-embeddings of graphs equipped with a bipartite dimer model [CLR1], also known as Coulomb gauges [KLRR], which simultaneously generalizes Tutte’s embeddings and the so-called s-embeddings of planar graphs carrying the Ising model. Shortly afterwards, it was understood that t-embeddings should be viewed as piecewise linear surfaces in the Minkowski space ${\mathbb{R}}^{2,2}$, called t-surfaces below, which provide a correct description of the complex structure arising in the small mesh size limit [chelkak-ramassamy-aztec, chelkak-semb, CLR2, berggren-nicoletti-russkikh-doubleaztec].
In this paper we apply the formalism of [CLR1] to Tutte’s embeddings and prove the convergence of discrete harmonic functions to solutions of the linearized Monge–Ampère equation under no regularity assumptions on the graphs in question. To see that linearized Monge–Ampère equations provide the right language for this problem, the reader may observe that in the degenerate case when the potential $\varphi$ is piecewise linear, the elliptic operator ${\mathcal{L}}_{\varphi}$ given by (1.9) generates the Brownian motion on the cable system (metric graph) of the corresponding Tutte’s embedding.
@@
To the best of our knowledge, recent literature on the subject was mostly focused on the very special case – so-called orthodiagonal tilings – which appears when Tutte’s embeddings of a graph and of its dual form non-self-intersecting quads. In this setup, the convergence to usual harmonic functions (that is, solutions of the equation $\Delta h=0$, which corresponds to the quadratic potential $\varphi(w)=\frac{1}{2}|w|^{2}$ in (1.9)) under no regularity assumptions on the graphs was obtained in [gurel-gurevich-jerison-nachmias]; see also references therein and [binder2024orthodiagonal, bou2024random]. However, this passage from Tutte’s embeddings to orthodiagonal ones is quite restrictive. In particular, it generally does not work if one starts with a given harmonic embedding $\Gamma_{\delta}$ and then superimposes $\Gamma_{\delta}$ with the harmonic embedding $\Gamma^{*}_{\delta}$ of the dual graph. (Even if edges of $\Gamma^{*}_{\delta}$ ‘stay close’ to the corresponding edges of $\Gamma_{\delta}$, the quads formed by them typically self-intersect.) In this paper we aim to provide a much more general framework that can be applied to the study of 2d lattice models on irregular weighted graphs given in advance and harmonically embedded into the plane, which avoids this constraint and also does not require that the potential $\varphi$ is quadratic.
@@
if $(v^{\ast}_{1}v^{\ast}_{2})$ is the dual edge to $(v_{1}v_{2})$ and $v^{\ast}_{1}$ is to the right from $(v_{1}v_{2})$. The harmonicity of ${\mathcal{H}}_{\delta}$ implies that the sum of these increments around each inner vertex of $\Gamma$ vanishes and hence ${\mathcal{H}}^{\ast}_{\delta}$ – which is called the dual Tutte’s embedding – is well defined up to a global additive constant. Let us emphasize that this dual embedding can have overlaps if $U$ is not convex; see Section 2 for details.
Let us now define piecewise constant maps $\Psi_{\delta}:U\to{\mathbb{C}}$ by setting
@@
(If $w\in{\mathbb{C}}$ is a vertex of $\Gamma_{\delta}$ or lies on an edge, we choose one of the incident faces arbitrarily.) A very important property of $\Psi_{\delta}$ is the existence of a piecewise linear function $\Phi_{\delta}:U\to{\mathbb{R}}$ such that
$\Psi_{\delta}=2\partial_{\bar{w}}\Phi_{\delta}=\partial_{x}\Phi_{\delta}+i\partial_{y}\Phi_{\delta}\quad\text{for all $w=x+iy$ lying on faces of $\Gamma_{\delta}$}.$
One can construct $\Phi_{\delta}$ explicitly by defining its values on vertices of $\Gamma_{\delta}$ via the identity
@@
and noting that $\Phi_{\delta}$ restricted to vertices of a face $v^{\ast}$ of $\Gamma_{\delta}$ coincides with a linear function with gradient ${\mathcal{H}}^{\ast}_{\delta}(v^{\ast})$. We call $\Phi_{\delta}$ the potential corresponding to $\Gamma_{\delta}$. Since ${\mathcal{H}}^{\ast}_{\delta}$ and hence $\Psi_{\delta}$ are defined up to a global additive constant, the potential $\Phi_{\delta}$ is defined by $\Gamma_{\delta}$ up to a linear function.
It follows from (1.1) and $c_{v_{1}v_{2}}\!>\!0$ that the piecewise linear function $\Phi_{\delta}$ is convex near each edge of $\Gamma_{\delta}$. Therefore, $\Phi_{\delta}$ is a convex function on each convex subset of $U$. The only property of potentials $\Phi_{\delta}$ that we use in our paper is their uniform convexity above scales $\delta\to 0$; see also Remark 1.3 below.
@@
We say that the potentials $(\Phi_{\delta})_{\delta\to 0}$ have property (CONV) on the domain $U\subset{\mathbb{C}}$ if there exist constants $\lambda>0$ and $C>0$ such that the inequalities
@@
The well-studied ‘nicest’ setup for discrete complex analysis beyond the square grid is that of rhombic lattices or isoradial graphs introduced by Duffin in [duffin-rhombic]; see also [chelkak-smirnov-isorad]. In this case one can easily deduce from (1.1) that the restriction of $\Phi_{\delta}$ onto the vertices of $\Gamma_{\delta}$ exactly equals the quadratic potential $\Phi_{\delta}(v)=\frac{1}{2}|{\mathcal{H}}_{\delta}(v)|^{2}$. In a more general setup of orthodiagonal tilings [gurel-gurevich-jerison-nachmias, binder2024orthodiagonal, bou2024random], one has $\Psi_{\delta}(w)=w+{O(\delta)}$ and hence $\Phi_{\delta}(w)=\frac{1}{2}|w|^{2}+{O(\delta|w|)}$, where $\delta\to 0$ denotes the size of the largest tile. As for the converse, let us emphasize that the assumption $\Psi_{\delta}(w)=w+{O(\delta)}$ does not put the analysis back to the setup of [gurel-gurevich-jerison-nachmias, binder2024orthodiagonal, bou2024random] even if $\delta$ is comparable with the lengths of all the edges of $\Gamma_{\delta}$ since the quads ${\mathcal{H}}_{\delta}(v_{1}){\mathcal{H}}^{\ast}_{\delta}(v^{\ast}_{1}){\mathcal{H}}_{\delta}(v_{2}){\mathcal{H}}^{\ast}_{\delta}(v^{\ast}_{2})$ may self-intersect. See also Remark 1.3 and Remark 1.4 below.
@@
It is not hard to see that property (CONV) for potentials $\Phi_{\delta}$ on $U$ implies property (LIP) for their gradients $\Psi_{\delta}$ on each subdomain $U^{\prime}\Subset U$ and vice versa; we refer the reader to Lemma 5.1 for the proof of this elementary fact. (It is sufficient to assume that $U^{\prime}$ lies in an $O(\delta)$-interior of $U$.) More importantly, property (LIP) is also equivalent to the $\kappa$-Lipschitzness, $\kappa=\kappa(\varkappa)<1$, of the so-called origami map in the terminology of [CLR1], which makes the techniques developed in [CLR1] applicable to the setup of this paper; see Section 2.2 below.
@@
give us a standard invariant measure for $X_{t}$ if one excludes the effect of the boundary of $\Gamma_{\delta}$ from the consideration. Let us formulate two important properties of these random walks.
@@
We say that a sequence $(\Gamma_{\delta})_{\delta\to 0}$ of Tutte’s embeddings has property (RW) on a domain $U^{\prime}\Subset U$ if there exists constants $c,C>0$ such that the following holds:
@@
Our first result shows the equivalence of the property (RW) of random walks on Tutte’s embeddings $\Gamma_{\delta}$ to (equivalent) ’geometric’ properties (CONV)/(LIP) of $\Gamma_{\delta}$ discussed above.
@@
The fact that (LIP) implies (RW) follows from the framework developed in [CLR1]; in particular, [CLR1, Proposition 6.4] gives the uniform ellipticity of the random walk $X_{t}$ above scale $\delta$. We recap this material in Section 2.3 and also sketch a self-contained proof of the ellipticity property in Remark 4.3. We prove the converse statement in Section 5; see Theorem 5.6.
@@ 1.3. Main convergence results
Assume now that potentials $(\Phi_{\delta})_{\delta\to 0}$ converge, uniformly on compacts, to a function $\varphi:U\to{\mathbb{R}}$. It follows from (1.4) and (1.5) that the limit $\varphi$ is uniformly convex and ${\mathrm{C}}^{1,1}$-smooth; the latter means that $\varphi$ is continuously differentiable and has uniformly Lipschitz gradient $\psi=2\varphi_{\bar{w}}=\varphi_{x}+i\varphi_{y}$. Let
(1.8) $A_{\varphi}\ =\ \begin{pmatrix}\varphi_{yy}&-\varphi_{xy}\\ -\varphi_{xy}&\varphi_{xx}\end{pmatrix}\ =\ \begin{pmatrix}\operatorname{\mathrm{Re}}(\psi_{w}-\psi_{\bar{w}})&\operatorname{\mathrm{Im}}(\psi_{w}-\psi_{\bar{w}})\\ -\operatorname{\mathrm{Im}}(\psi_{w}+\psi_{\bar{w}})&\operatorname{\mathrm{Re}}(\psi_{w}+\psi_{\bar{w}})\,.\end{pmatrix}$
Note that $A_{\varphi}$ is defined almost everywhere due to Rademacher’s theorem and is uniformly elliptic, i.e., $\lambda\leq A_{\varphi}\leq\lambda^{-1}$ due to (1.4). Consider the differential operator
@@
The equation ${\mathcal{L}}_{\varphi}h=0$ is known under the name linearized Monge–Ampère equation [caffarelli-gutierrez-1997-lin-MA]. Even though the matrix $A_{\varphi}$ is only measurable, the fact that the operator ${\mathcal{L}}_{\varphi}$ can also be written in the divergence form allows one to speak about weak solutions, that is, functions $h\in W^{1,2}_{\mathrm{loc}}(\Omega)$ such that $\int_{\Omega}(A_{\varphi}\nabla h\cdot\nabla\phi)(w)d^{2}w=0$ for all smooth test functions $\phi\in{\mathrm{C}}_{0}^{\infty}(\Omega)$.
Another important feature coming from the divergence expression for ${\mathcal{L}}_{\varphi}$ is the notion of a harmonic conjugate. Define $\ast=\left(\begin{smallmatrix}0&-1\\ 1&0\end{smallmatrix}\right)$ and note that for every weak solution $h$ there is a function $h^{\ast}\in W_{\mathrm{loc}}^{1,2}(\Omega)$ satisfying $\nabla h^{\ast}=\ast A_{\varphi}\nabla h$. Such $h^{\ast}$ is called an $A_{\varphi}$-harmonic conjugate of $h$; see [AstalaBook, Chapter 16] and Section 3 for more details. Note that $h^{\ast}$ is a weak solution of the ‘dual’ equation
@@
The standard theory of elliptic equations (see [kenig2000new] and references therein) guaranties that weak solutions $h,h^{*}$ are continuous in $\Omega$. Moreover, the Dirichlet problem ${\mathcal{L}}_{\varphi}h=0$ in $\Omega$, $h|_{\partial\Omega}=g$ has a unique solution $h\in W^{1,2}_{\mathrm{loc}}(\Omega)\cap{\mathrm{C}}(\overline{\Omega})$ for each continuous function $g:\partial\Omega\to{\mathbb{R}}$ [littman-stampacchia-weinberger-1963-regular-points, kenig2000new].
@@
Instead of assuming that $(\Gamma_{\delta})_{\delta\to 0}$ have property (CONV), one can start with an arbitrary sequence of Tutte’s embeddings $\Gamma^{(n)}$ such that the potentials $\Phi^{(n)}$ converge to a uniformly convex function $\varphi:U\to{\mathbb{R}}$. This implies the existence of parameters $\delta^{(n)}\to 0$ such that $\Phi^{(n)}$ have property (CONV) with twice smaller $\lambda$ starting from the scale $\delta^{(n)}$, and one can rename $\Phi^{(n)}$ by $\Phi_{\delta^{(n)}}$.
@@
We prove Theorem 2 and Theorem 3 in Section 4 first under an additional smoothness assumption $\varphi\in{\mathrm{C}}^{3}$ in Section 4.3 and then in the general case $\varphi\in{\mathrm{C}}^{1,1}$. A simplification comes from the fact that for $\varphi\in{\mathrm{C}}^{3}$ each ‘very weak’ solution of the equation ${\mathcal{L}}_{\varphi}h=0$, i.e., a function $h$ such that $\int_{\Omega}h(w)[{\mathcal{L}}_{\varphi}\phi](w)d^{2}w=0$ for all $\phi\in{\mathrm{C}}^{\infty}_{0}(\Omega)$, is automatically a weak solution. In the general case $\varphi\in{\mathrm{C}}^{1,1}$, we rely upon the a priori regularity theory developed in [CLR1] and an idea somewhat similar to the one used in [mahfouf-crossing-estimates] in the Ising model context in order to show that each discrete harmonic function on $\Gamma_{\delta}$ can be locally approximated by discrete harmonic functions on the same graph with bounded gradients. However, we do not know whether or not all discrete harmonic functions on $\Gamma_{\delta}$ have uniformly bounded gradients. In the positive direction, we prove that (a) the convergence in Theorem 2 and 3 also holds for $\nabla H_{\delta}$ and $\nabla G_{\Omega}^{\delta}(\cdot,v_{0})$ if the graphs $(\Gamma_{\delta})_{\delta\to 0}$ satisfy a very mild additional regularity assumption (EXP-FAT), and (b) the same convergence always holds for the ‘gradients defined on a mesoscopic scale’. See Section 4.4 and Section 4.5 for details.
@@
In the special setup of orthodiagonal tilings, Theorem 2 has been obtained in [gurel-gurevich-jerison-nachmias]; see also references therein and [bou2024random, Theorem B] for an alternative proof. Moreover, in [bou2024random, Theorem A] the authors show that in this case [bou2024random, Theorem B] implies the convergence of trajectories of the random walks on $\Gamma_{\delta}$ to trajectories of the Brownian motion, up to a possible time-reparametrization. We believe that Theorem 2 supplemented by Theorem 3 is sufficient to prove the convergence of random walks $X_{t}$ on $\Gamma_{\delta}$ to the driftless diffusion defined by the elliptic operator ${\mathcal{L}}_{\varphi}$ together with the time-parametrizations. However, we do not explore this question in our paper.
@@
In applications of convergence results to 2d lattice models, it is often important to drop the smoothness assumption on $\partial\Omega$ and to use the Carathéodory topology for the convergence $\Omega_{\delta}\to\Omega$ as $\delta\to 0$. The reason for such a generalization stems from the fact that it allows to use a simple compactness argument in order to deduce that the convergence, e.g., of functions $H_{\delta}$ to $h$ in Theorem 2 is uniform with respect to the domain $\Omega$; e.g., see a discussion in [chelkak-smirnov-isorad, Section 3.2]. In some setups it is also important to consider harmonic measures of boundary arcs or fractal interfaces instead of solutions of Dirichlet problems with continuous boundary data $g$. Such generalizations of Theorem 2 and Theorem 3 can be deduced from the uniform ellipticity property (RW) of random walks on $\Gamma_{\delta}$, which implies the so-called weak-Beurling estimate similarly to [chelkak-smirnov-isorad].

### v2

@@ 1.1. Motivation
Convergence of discrete harmonic functions in the small mesh size limit is a classical topic dating back to the first works on numerical methods [33, 17]. In two dimensions, this question is intrinsically related to the notion of discrete holomorphic functions and their convergence, with important contributions made in the middle of the last century by Ferrand [30, Chapter V] who greatly developed this theory on the square grid and Duffin [19] who, in particular, suggested a generalization to the so-called rhombic tilings of the complex plane. From the combinatorial perspective, two related classical setups for which the notions of discrete holomorphic and harmonic functions make perfect sense are Brooks–Stone–Smith–Tutte’s square (or, more generally, rectangular) tilings [8] and Tutte’s barycentric, or harmonic, embeddings [45]. One of the sources of the modern interest to discrete complex analysis is the 2d statistical physics; e.g., see [43]. In particular, the framework of rhombic lattices was important for the study of scaling limits of free fermionic models with Baxter’s Z-invariant weights; e.g., see [37, 28, 14, 7].
@@
Recent advances of the probabilistic approach to the Liouville Quantum Gravity (LQG) in 2d (e.g., see [25] and references therein) motivate a development of discrete complex analysis techniques for possibly very irregular planar grids that appear as embeddings of given (random) planar graphs into the plane. In particular, applications of Tutte’s embeddings and square tilings in the LQG context recently appeared in [26, 4]. Going in this direction, the last three authors of this paper suggested a general framework of the so-called t-embeddings of graphs equipped with a bipartite dimer model [11], also known as Coulomb gauges [29], which simultaneously generalizes Tutte’s embeddings and the so-called s-embeddings of planar graphs carrying the Ising model. Shortly afterwards, it was understood that t-embeddings should be viewed as piecewise linear surfaces in the Minkowski space ${\mathbb{R}}^{2,2}$, called t-surfaces below, which provide a correct description of the complex structure arising in the small mesh size limit [12, 16, 10, 3].
In this paper we apply the formalism of [11] to Tutte’s embeddings and prove the convergence of discrete harmonic functions to solutions of the linearized Monge–Ampère equation under no regularity assumptions on the graphs in question. The potential $\varphi$ of the equation is obtained as a limit of piecewise linear potentials associated with the embeddings via the Maxwell–Cremona correspondence (see Section 1.2). We emphasize that in general $\varphi$ can be an arbitrary uniformly convex function (see (1.4)), and so the limits of discrete harmonic functions might not be harmonic in the usual sense even after a non-trivial change of coordinates. If such coordinates exist, they must be isothermal with respect to the Lorentz metric on the corresponding t-surface, see Section 1.4.
@@
To the best of our knowledge, recent literature on the subject was mostly focused on the very special case – so-called orthodiagonal tilings – which appears when Tutte’s embeddings of a graph and of its dual form non-self-intersecting quads. In this setup, the convergence to usual harmonic functions (that is, solutions of the equation $\Delta h=0$, which corresponds to the quadratic potential $\varphi(w)=\frac{1}{2}|w|^{2}$ in (1.9)) under no regularity assumptions on the graphs was obtained in [23]; see also references therein and [5, 6]. However, this passage from Tutte’s embeddings to orthodiagonal ones is quite restrictive. In particular, it generally does not work if one starts with a given harmonic embedding $\Gamma_{\delta}$ and then superimposes $\Gamma_{\delta}$ with the harmonic embedding $\Gamma^{*}_{\delta}$ of the dual graph. (Even if edges of $\Gamma^{*}_{\delta}$ ‘stay close’ to the corresponding edges of $\Gamma_{\delta}$, the quads formed by them typically self-intersect.) In this paper we aim to provide a much more general framework that can be applied to the study of 2d lattice models on irregular weighted graphs given in advance and harmonically embedded into the plane, which avoids this constraint and also does not require that the potential $\varphi$ is quadratic.
@@
if $(v^{\ast}_{1}v^{\ast}_{2})$ is the dual edge to $(v_{1}v_{2})$ and $v^{\ast}_{1}$ is to the right from $(v_{1}v_{2})$. The harmonicity of ${\mathcal{H}}_{\delta}$ implies that the sum of these increments around each inner vertex of $\Gamma$ vanishes and hence ${\mathcal{H}}^{\ast}_{\delta}$ – which is called the dual Tutte’s embedding – is well defined up to a global additive constant. Let us emphasize that this dual embedding can have overlaps if the boundary of ${\mathcal{H}}_{\delta}(\Gamma_{\delta})$ is not convex; see Section 2 for details.
A crucial role for our results is played by the Maxwell–Cremona lift of ${\mathcal{H}}_{\delta}$. To define this lift we first define the piecewise constant map $\Psi_{\delta}:U\to{\mathbb{C}}$ by setting
@@
(If $w\in{\mathbb{C}}$ is a vertex of $\Gamma_{\delta}$ or lies on an edge, we choose one of the incident faces arbitrarily.) We then define a piecewise linear continuous function $\Phi_{\delta}:U\to{\mathbb{R}}$ by requiring that
$\Psi_{\delta}=2\partial_{\bar{w}}\Phi_{\delta}=\partial_{x}\Phi_{\delta}+i\partial_{y}\Phi_{\delta}\quad\text{for all $w=x+iy$ lying on faces of $\Gamma_{\delta}$}$
or, equivalently, that $\Phi_{\delta}$ is linear on faces of ${\mathcal{H}}_{\delta}(\Gamma_{\delta})$ and
@@
for every two incident vertices $v_{1},v_{2}$ of $\Gamma$, where $v_{1}^{\ast}$ and $v_{2}^{\ast}$ are the faces incident to $v_{1}v_{2}$. The classic Maxwell theorem [35, 36, 18] asserts that $\Phi_{\delta}$ is well-defined if and only if the collection of the (not necessary positive) weights $(c_{vv^{\prime}})$ is such that the map ${\mathcal{H}}_{\delta}$ is a harmonic embedding and ${\mathcal{H}}^{\ast}_{\delta}$ satisfies (1.1). If the 1-skeleton of ${\mathcal{H}}_{\delta}(\Gamma_{\delta})$ is fixed, this defines a correspondence between polygonal lifts of ${\mathcal{H}}_{\delta}(\Gamma_{\delta})$ to ${\mathbb{R}}^{3}$ and such edge-weights (or self-stresses), called the Maxwell–Cremona correspondence [39, 40].
It follows from (1.1) that $c_{v_{1}v_{2}}\!>\!0$ if and only if the piecewise linear function $\Phi_{\delta}$ is convex near each edge of $\Gamma_{\delta}$. In this case $\Phi_{\delta}$ is a convex function on each convex subset of $U$ and we call it the Maxwell–Cremona potential of ${\mathcal{H}}_{\delta}$. The only property of potentials $\Phi_{\delta}$ that we use in our paper is their uniform convexity above scales $\delta\to 0$. We emphasize that $\delta$ can be much larger then the size of any face of ${\mathcal{H}}_{\delta}(\Gamma_{\delta})$; see also Remark 1.3 below.
@@
We say that the potentials $(\Phi_{\delta})_{\delta\to 0}$ have the property (CONV) on the domain $U\subset{\mathbb{C}}$ if there exist constants $\lambda>0$ and $C>0$ such that the inequalities
@@
The well-studied ‘nicest’ setup for discrete complex analysis beyond the square grid is that of rhombic lattices or isoradial graphs introduced by Duffin in [19]; see also [13]. In this case one can easily deduce from (1.1) that the restriction of $\Phi_{\delta}$ onto the vertices of $\Gamma_{\delta}$ exactly equals the quadratic potential $\Phi_{\delta}(v)=\frac{1}{2}|{\mathcal{H}}_{\delta}(v)|^{2}$. In a more general setup of orthodiagonal tilings [23, 5, 6], one has $\Psi_{\delta}(w)=w+{O(\delta)}$ and hence $\Phi_{\delta}(w)=\frac{1}{2}|w|^{2}+{O(\delta|w|)}$, where $\delta\to 0$ denotes the size of the largest tile. As for the converse, let us emphasize that the assumption $\Psi_{\delta}(w)=w+{O(\delta)}$ does not put the analysis back to the setup of [23, 5, 6] even if $\delta$ is comparable with the lengths of all the edges of $\Gamma_{\delta}$ since the quads ${\mathcal{H}}_{\delta}(v_{1}){\mathcal{H}}^{\ast}_{\delta}(v^{\ast}_{1}){\mathcal{H}}_{\delta}(v_{2}){\mathcal{H}}^{\ast}_{\delta}(v^{\ast}_{2})$ may self-intersect. See also Remark 1.3 and Remark 1.4 below.
@@
It is not hard to see that property (CONV) for potentials $\Phi_{\delta}$ on $U$ implies property (LIP) for their gradients $\Psi_{\delta}$ on each subdomain $U^{\prime}\Subset U$ and vice versa; we refer the reader to Lemma 5.1 for the proof of this elementary fact. (It is sufficient to assume that $U^{\prime}$ lies in an $O(\delta)$-interior of $U$.) More importantly, property (LIP) is also equivalent to the $\kappa$-Lipschitzness, $\kappa=\kappa(\varkappa)<1$, of the so-called origami map in the terminology of [11], which makes the techniques developed in [11] applicable to the setup of this paper; see Section 2.2 below.
@@
give us a standard invariant measure for $X_{t}$ if one excludes the effect of the boundary of $\Gamma_{\delta}$ from the consideration. Let us formulate an important property of these random walks.
@@
We say that a sequence $(\Gamma_{\delta})_{\delta\to 0}$ of Tutte’s embeddings has property (RW) on a domain $U^{\prime}\Subset U$ if there exist constants $c,C>0$ such that the following holds:
@@
Our first result shows the equivalence of the property (RW) of random walks on Tutte’s embeddings $\Gamma_{\delta}$ to the (equivalent) ‘geometric’ properties (CONV)/(LIP) of $\Gamma_{\delta}$ discussed above.
@@
The fact that (LIP) implies (RW) follows from the framework developed in [11]; in particular, [11, Proposition 6.4] gives the uniform ellipticity of the random walk $X_{t}$ above scale $\delta$. We recap this material in Section 2.3 and also sketch a self-contained proof of the ellipticity property in Remark 4.3. We prove the converse statement in Section 5; see Theorem 5.6.
@@ 1.3. Main convergence results
Assume now that Maxwell–Cremona potentials $(\Phi_{\delta})_{\delta\to 0}$ converge, uniformly on compacts, to a function $\varphi:U\to{\mathbb{R}}$. It follows from (1.4) and (1.5) that the limit $\varphi$ is uniformly convex and ${\mathrm{C}}^{1,1}$-smooth; the latter means that $\varphi$ is continuously differentiable and has uniformly Lipschitz gradient $\psi=2\varphi_{\bar{w}}=\varphi_{x}+i\varphi_{y}$. Let
(1.8) $A_{\varphi}\ =\ \begin{pmatrix}\varphi_{yy}&-\varphi_{xy}\\ -\varphi_{xy}&\varphi_{xx}\end{pmatrix}\ =\ \begin{pmatrix}\operatorname{\mathrm{Re}}(\psi_{w}-\psi_{\bar{w}})&\operatorname{\mathrm{Im}}(\psi_{w}-\psi_{\bar{w}})\\ -\operatorname{\mathrm{Im}}(\psi_{w}+\psi_{\bar{w}})&\operatorname{\mathrm{Re}}(\psi_{w}+\psi_{\bar{w}})\,\end{pmatrix}$.
Note that $A_{\varphi}$ is defined almost everywhere due to Rademacher’s theorem and is uniformly elliptic, i.e. $\lambda\leq A_{\varphi}\leq\lambda^{-1}$ a.e., due to (1.4). Consider the differential operator
@@
The equation ${\mathcal{L}}_{\varphi}h=0$ is known under the name linearized Monge–Ampère equation [9]. Even though the matrix $A_{\varphi}$ is only measurable, the fact that the operator ${\mathcal{L}}_{\varphi}$ can also be written in the divergence form allows one to speak about weak solutions, that is, functions $h\in W^{1,2}_{\mathrm{loc}}(\Omega)$ such that $\int_{\Omega}(A_{\varphi}\nabla h\cdot\nabla\phi)(w)d^{2}w=0$ for all smooth test functions $\phi\in{\mathrm{C}}_{0}^{\infty}(\Omega)$.
Another important feature coming from the divergence expression for ${\mathcal{L}}_{\varphi}$ is the notion of a harmonic conjugate. Define $\ast=\left(\begin{smallmatrix}0&-1\\ 1&0\end{smallmatrix}\right)$ and note that for every weak solution $h$ there is a function $h^{\ast}\in W_{\mathrm{loc}}^{1,2}(\Omega)$ satisfying $\nabla h^{\ast}=\ast A_{\varphi}\nabla h$. Such $h^{\ast}$ is called an $A_{\varphi}$-harmonic conjugate of $h$; see [1, Chapter 16] and Section 3 for more details. Note that $h^{\ast}$ is a weak solution of the ‘dual’ equation
@@
The standard results from the theory of elliptic equations (see [27] and references therein) guaranty that weak solutions $h,h^{*}$ are continuous in $\Omega$. Moreover, the Dirichlet problem ${\mathcal{L}}_{\varphi}h=0$ in $\Omega$, $h|_{\partial\Omega}=g$, has a unique solution $h\in W^{1,2}_{\mathrm{loc}}(\Omega)\cap{\mathrm{C}}(\overline{\Omega})$ for each continuous function $g:\partial\Omega\to{\mathbb{R}}$ [31, 27].
@@
Instead of assuming that $(\Gamma_{\delta})_{\delta\to 0}$ have property (CONV), one can start with an arbitrary sequence of Tutte’s embeddings $\Gamma^{(n)}$ such that the potentials $\Phi^{(n)}$ converge uniformly to a uniformly convex function $\varphi:U\to{\mathbb{R}}$. This implies the existence of parameters $\delta^{(n)}\to 0$ such that $\Phi^{(n)}$ have property (CONV) with twice smaller $\lambda$ starting from the scale $\delta^{(n)}$, and one can rename $\Phi^{(n)}$ by $\Phi_{\delta^{(n)}}$.
@@
We prove Theorem 2 and Theorem 3 in Section 4: first under the additional smoothness assumption $\varphi\in{\mathrm{C}}^{3}$ in Section 4.3 and then in the general case $\varphi\in{\mathrm{C}}^{1,1}$ in Section 4.5. The principle difference between the general case and $\varphi\in{\mathrm{C}}^{3}$ is the regularity lemma asserting that, under the assumption $\varphi\in{\mathrm{C}}^{3}$, every ‘very weak’ solution of the equation ${\mathcal{L}}_{\varphi}h=0$, i.e., a function $h$ such that $\int_{\Omega}h(w)[{\mathcal{L}}_{\varphi}\phi](w)d^{2}w=0$ for all $\phi\in{\mathrm{C}}^{\infty}_{0}(\Omega)$, is automatically a weak solution (see Lemma 3.2). In the general case $\varphi\in{\mathrm{C}}^{1,1}$ such a statement requires the existence of the gradient of $h$ to be known a priori. To overcome this difficulty we choose to rely upon the a priori regularity theory developed in [11] and an idea somewhat similar to the one used in [34] in the Ising model context in order to show that each discrete harmonic function on $\Gamma_{\delta}$ can be locally approximated by discrete harmonic functions on the same graph with bounded gradients. This allows us to control the gradients of limits of discrete harmonic functions. Moreover, as a byproduct of this approach we show that the convergence in Theorem 2 and 3 holds also for the ‘gradients defined on a mesoscopic scale’. While we do not know whether or not the actual gradients of discrete harmonic functions can be bounded in the general setting, we prove that the convergence in Theorem 2 and 3 also holds for $\nabla H_{\delta}$ and $\nabla G_{\Omega}^{\delta}(\cdot,v_{0})$ if the graphs $(\Gamma_{\delta})_{\delta\to 0}$ satisfy a very mild additional regularity assumption (EXP-FAT). See Section 4.4 and Section 4.5 for details.
@@
In the special setup of orthodiagonal tilings, Theorem 2 has been obtained in [23]; see also references therein and [6, Theorem B] for an alternative proof. Moreover, in [6, Theorem A] the authors show that in this case [6, Theorem B] implies the convergence of trajectories of the random walks on $\Gamma_{\delta}$ to trajectories of the Brownian motion, up to a possible time-reparametrization. We believe that Theorem 2 supplemented by Theorem 3 is sufficient to prove the convergence of random walks $X_{t}$ on $\Gamma_{\delta}$ to the driftless diffusion defined by the elliptic operator ${\mathcal{L}}_{\varphi}$ together with the time-parametrizations. However, we do not explore this question in our paper.
@@
Although $\Phi_{\delta}$ is not twice differentiable, one can still define the quadratic form of the operator ${\mathcal{L}}_{\Phi_{\delta}}$ acting, say, on pairs of smooth compactly supported functions. A direct examination shows that in this case the elliptic operator ${\mathcal{L}}_{\Phi_{\delta}}$ generates the Brownian motion on the cable system (metric graph) of the corresponding Tutte’s embedding. This explains why linearized Monge–Ampère equations provide the right language for the problem.
Remark 1.6.
In applications of convergence results to 2d lattice models, it is often important to drop the smoothness assumption on $\partial\Omega$ and to use the Carathéodory topology for the convergence $\Omega_{\delta}\to\Omega$ as $\delta\to 0$. The reason for such a generalization stems from the fact that it allows to use a simple compactness argument in order to deduce that the convergence, e.g., of functions $H_{\delta}$ to $h$ in Theorem 2 is uniform with respect to the domain $\Omega$; e.g., see a discussion in [13, Section 3.2]. In some setups it is also important to consider harmonic measures of boundary arcs or fractal interfaces instead of solutions of Dirichlet problems with continuous boundary data $g$. Such generalizations of Theorem 2 and Theorem 3 can be deduced from the uniform ellipticity property (RW) of random walks on $\Gamma_{\delta}$, which implies the so-called weak-Beurling estimate similarly to [13].